# Thought Anchors: Which LLM Reasoning Steps Matter?

### v2

@@ Appendix A Sentence taxonomy
Building on top of the framework presented by [31], we developed a taxonomy consisting of eight distinct sentence categories that capture reasoning functions in mathematical problem-solving. Each category represents a specific cognitive operation. The functions and examples for each category are given in Table 1. Notably, the uncertainty management category includes backtracking sentences.
@@ D.2 Attention aggregation
The R1-Distill-Qwen-14B and R1-Distill-Llama-8B receiver heads displayed consistent patterns related to sentence types, such that plan generation, uncertainty management, and self checking sentences received heightened attention; although visually, the differences to fact retrieval and active computation may be less prominent, paired t-tests (paired with respect to a given response) showed that plan generation, uncertainty management, and self checking always significantly surpassed fact retrieval and active computation (six paired t-tests: $p$s .$\leq.01$).
@@ Appendix E Embeddings model
We used all-MiniLM-L6-v2 with a maximum sequence length of $256$ tokens and a hidden dimension of $384$ as our sentence embeddings model from the sentence-transformers [27] library. We picked a cosine similarity threshold of $0.8$, which is the median similarity value between all sentence removed (i.e., original sentence) and sentence resampled pairs in our dataset.
@@ G.1 Counterfactual versus resampling importance
However, the counterfactual importance metric can yield high-variance estimates when the number of semantically divergent resampled sentences is low (e.g., $<10$), as the conditional probability estimates become less reliable with limited data. Alongside the limitations discussed in Section 8, this represents another constraint of our approach that future work should investigate further.
@@ G.2 Sentence-to-sentence importance
  • 8-FR, 9-PG, 12-PG, 14-PG $\rightarrow$ 32-UM. The uncertainty management in sentence 32 (“However, I need to check if 440-449 is fully included.”) receives sizeable positive influence from several earlier sentences: 8-FR ($+0.11$), 9-PG ($+0.06$), 12-PG ($+0.12$), 14-PG ($+0.10$). This forms the dense horizontal band at row index 32.
  • 39-RC ↛$\not\rightarrow$ 83-UM. The result consolidation in sentence 32 (“Now, summing up all the 4’s: - Units: 48 - Tens: 50 - Hundreds: 89. Total 4’s = 48 + 50 + 89 = 187.”) decreases the likelihood (i.e., inhibits) of 83-UM (“Wait, but just to be thorough, let me check the hundreds place for 4’s again.”) by $0.22$.
  • 52-AC ↛$\not\rightarrow$ 65-SC. The computation in sentence 52 (“The first four blocks 80-89, 180-189, 280-289, 380-389 each contribute 10 eights in the tens place.”) decreases the likelihood of 65-SC (“Let me go through each step again to make sure I didn’t make a mistake.”) by $0.16$.
  • 63-AC $\rightarrow$ 64-UM, 65-SC, 69-SC, 75-SC, 83-UM, 86-SC. The computation in sentence 63 (“So, the difference is 187 – 98 = 89.”) propagates forward, increasing the likelihood of 64-UM ($+0.24$), 65-SC ($+0.17$), 69-SC ($+0.16$), 75-SC ($+0.28$), 83-UM ($0.23$), and 86-SC ($0.16$). This forms the dense vertical band originating from column index 63.
  • 64-UM $\rightarrow$ 65-SC, 69-SC, 75-SC, 83-UM, 86-SC. The uncertainty management in sentence 64 (“Wait, that seems quite a large difference.”) further amplifies the same downstream block: 65-SC ($+0.32$), 69-SC ($+0.25$), 75-SC ($+0.26$), 83-UM ($0.25$), and 86-SC ($0.25$).
@@
Figure 16: Sentence-to-sentence importance matrix for the 32 most important sentences in problem #2236 (incorrect), selected based on total outgoing and incoming importance. Each cell $(i,j)$ shows the causal importance of sentence $i$ on sentence $j$, calculated as the difference in the probability sentence $j$ semantically occurs ($>0.8$ cosine similarity) when sentence $i$ is present versus resampled.
@@ Appendix I Reasoning versus base model differences in receiver heads
To test the causal hypothesis that the receiver heads identified in Section 4 are functionally important for reasoning, we performed an experiment ablating receiver heads and evaluating how this impact’s model accuracy. This intervention is designed to measure the direct impact of removing these heads on task performance and to evaluate the possibility that they may be more important than typical heads.
@@
[…]
@@
A sentence’s receiver-head score predicts its impact on downstream sentences per the resampling method. Specifically, our resampling approach (Section 3) produces a matrix describing each sentence’s counterfactual importance on each future sentence. Based on said matrix, we can compute a sentence-level measure analogous to the receiver head score by taking the average value below the matrix diagonal; as with the receiver head score, we omit the four entries immediately below the diagonal to avoid simply capturing proximal effects and focus on farther downstream impacts; we take the absolute value of the resampling matrix to produce an estimate of how a given sentence may both upregulate or downregulate future sentences.
[…]
@@
Below is a transcript of the model’s response used for the case study (Section 6). The first 74 sentences following <think> are shown. After this point, the model converged on 100% resampling accuracy for all subsequent sentences.
@@
    (fact_retrieval): Starting from the right, the digits are $16^{0}$,$16^{1}$,$16^{2}$,$16^{3}$,$16^{4}$.

### v3

@@ Appendix A Sentence taxonomy
Building on top of the framework presented by [Venhoff et al., 2025], we developed a taxonomy consisting of eight distinct sentence categories that capture reasoning functions in mathematical problem-solving. Each category represents a specific cognitive operation. The functions and examples for each category are given in Table 1. Notably, the uncertainty management category includes backtracking sentences.
@@ D.2 Attention aggregation
The R1-Distill-Qwen-14B and R1-Distill-Llama-8B receiver heads displayed consistent patterns related to sentence types, such that plan generation, uncertainty management, and self checking sentences received heightened attention; although visually, the differences to fact retrieval and active computation may be less prominent, paired t-tests (paired with respect to a given response) showed that plan generation, uncertainty management, and self checking always significantly surpassed fact retrieval and active computation (six paired t-tests: $p$s $\leq.01$).
@@ Appendix E Embeddings model
We used all-MiniLM-L6-v2 with a maximum sequence length of $256$ tokens and a hidden dimension of $384$ as our sentence embeddings model from the sentence-transformers [Reimers and Gurevych, 2019] library. We picked a cosine similarity threshold of $0.8$, which is the median similarity value between all sentence removed (i.e., original sentence) and sentence resampled pairs in our dataset.
@@ G.1 Counterfactual versus resampling importance
However, the counterfactual importance metric can yield high-variance estimates when the number of semantically divergent resampled sentences is low (e.g., <10<10), as the conditional probability estimates become less reliable with limited data. Alongside the limitations discussed in Section 8, this represents another constraint of our approach that future work should investigate further.
@@ G.2 Sentence-to-sentence importance
  • 8-FR, 9-PG, 12-PG, 14-PG $\rightarrow$ 32-UM. The uncertainty management in sentence 32 (“However, I need to check if 440-449 is fully included.”) receives sizeable positive influence from several earlier sentences: 8-FR (+0.11+0.11), 9-PG (+0.06+0.06), 12-PG (+0.12+0.12), 14-PG (+0.10+0.10). This forms the dense horizontal band at row index 32.
  • 39-RC $\not\rightarrow$ 83-UM. The result consolidation in sentence 32 (“Now, summing up all the 4’s: - Units: 48 - Tens: 50 - Hundreds: 89. Total 4’s = 48 + 50 + 89 = 187.”) decreases the likelihood (i.e., inhibits) of 83-UM (“Wait, but just to be thorough, let me check the hundreds place for 4’s again.”) by $0.22$.
  • 52-AC $\not\rightarrow$ 65-SC. The computation in sentence 52 (“The first four blocks 80-89, 180-189, 280-289, 380-389 each contribute 10 eights in the tens place.”) decreases the likelihood of 65-SC (“Let me go through each step again to make sure I didn’t make a mistake.”) by $0.16$.
  • 63-AC $\rightarrow$ 64-UM, 65-SC, 69-SC, 75-SC, 83-UM, 86-SC. The computation in sentence 63 (“So, the difference is 187 – 98 = 89.”) propagates forward, increasing the likelihood of 64-UM (+0.24+0.24), 65-SC (+0.17+0.17), 69-SC (+0.16+0.16), 75-SC (+0.28+0.28), 83-UM ($0.23$), and 86-SC ($0.16$). This forms the dense vertical band originating from column index 63.
  • 64-UM $\rightarrow$ 65-SC, 69-SC, 75-SC, 83-UM, 86-SC. The uncertainty management in sentence 64 (“Wait, that seems quite a large difference.”) further amplifies the same downstream block: 65-SC (+0.32+0.32), 69-SC (+0.25+0.25), 75-SC (+0.26+0.26), 83-UM ($0.25$), and 86-SC ($0.25$).
@@
Figure 16: Sentence-to-sentence importance matrix for the 32 most important sentences in problem #2236 (incorrect), selected based on total outgoing and incoming importance. Each cell $(i,j)$ shows the causal importance of sentence $i$ on sentence $j$, calculated as the difference in the probability sentence $j$ semantically occurs (>0.8>0.8 cosine similarity) when sentence $i$ is present versus resampled.
@@ Appendix I Reasoning versus base model differences in receiver heads
To test the causal hypothesis that the receiver heads identified in Section˜4 are functionally important for reasoning, we performed an experiment ablating receiver heads and evaluating how this impact’s model accuracy. This intervention is designed to measure the direct impact of removing these heads on task performance and to evaluate the possibility that they may be more important than typical heads.
@@
The importance of receiver heads emerges when a large number of heads are ablated. When ablating 512 heads (over a quarter of the model’s 1920 heads), targeting receiver heads caused performance to fall to 28% accuracy. Removing the same number of control heads resulted in a less severe drop to 37% accuracy. There is a significant difference between these percentages ($t[31]=2.55,p=.02$), suggesting receiver heads are more critical for reasoning than other heads.
As mentioned, this analysis treats responses as incorrect if they do not produce a final answer by 16,384 tokens. If the analysis is changed to instead simply omit those responses entirely from the analysis, there remains a significant difference in accuracy when ablating top-512 receiver heads (29% accuracy) versus random non-receiver heads (39% accuracy) ($t[31]=2.66,p=.02)$. Hence, regardless of whether non-completed responses are marked as incorrect or ignored, ablating receiver heads is found to exert a larger impact on model accuracy than ablating random non-receiver heads.
@@
A sentence’s receiver-head score predicts its impact on downstream sentences per the resampling method. Specifically, our resampling approach (Section˜3) produces a matrix describing each sentence’s counterfactual importance on each future sentence. Based on said matrix, we can compute a sentence-level measure analogous to the receiver head score by taking the average value below the matrix diagonal; as with the receiver head score, we omit the four entries immediately below the diagonal to avoid simply capturing proximal effects and focus on farther downstream impacts; we take the absolute value of the resampling matrix to produce an estimate of how a given sentence may both upregulate or downregulate future sentences.
We correlated sentence (top-64-head) receiver scores with their resampling scores calculated here - correlated separately for each response - yielding a positive association (mean Spearman $\rho=.22$; 95% CI: [.17, .30]). Although the correlation is nominally weak/moderate, this is a consistent trend, and 18 of the 20 responses tested displayed a positive association ($\rho>0$). A sentence’s receiver-score additionally correlates with its downstream effect as computed by the attention suppression measure introduced in Section˜5 (mean >4 below the diagonal) (mean Spearman $\rho=.19$; 95% CI: [.10, .28]). Although attention weights alone are not a causal measure, they are computationally cheap to examine, requiring no post hoc forward pass, unlike the other two methods. Grounded in these other techniques, the receiver-head scores may hence prove useful by virtue of their limited cost.
@@
Below is a transcript of the model’s response used for the case study (Section˜6). The first 74 sentences following <think> are shown. After this point, the model converged on 100% resampling accuracy for all subsequent sentences.
@@
    (fact_retrieval): Starting from the right, the digits are $16^{0}$$16^{1}$$16^{2}$$16^{3}$$16^{4}$.